OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of 1/3 and passes through the point (-6, 0).

OpenStudy (anonymous):

y = x/3 + 2

OpenStudy (help!!!!):

use point slope equation \[y-y_{1}=m(x-x _{1})\]

OpenStudy (help!!!!):

Then write it in y-intercept form

OpenStudy (anonymous):

the answers choses are a: \[y-0=\frac{ 1 }{ 3}\left( x+6 \right)\] b:\[y-0=\frac{ 1 }{ 3 }\left( x-6 \right)\] c.\[y+6=\frac{ 1 }{ 3 }\left( x-0 \right)\] d. \[y-6=\frac{ 1 }{ 3 }\left( x-0 \right)\]